# Spectrum and low-energy gap in triangular quantum spin liquid NaYbSe2

### v1

@@ Abstract
We report neutron scattering, pressure-dependent AC calorimetry, and AC magnetic susceptibility measurements of triangular lattice NaYbSe2. We observe a continuum of scattering, which is reproduced by matrix product simulations, and no phase transition is detected in any bulk measurements. Comparison to heat capacity simulations suggest the material is within the Heisenberg spin liquid phase. AC Susceptibility shows a significant 23 mK downturn, indicating a gap in the magnetic spectrum. The combination of a gap with no detectable magnetic order, comparison to theoretical models, and comparison to other $A$YbSe2 compounds all strongly indicate NaYbSe2 is within the quantum spin liquid phase. The gap also allows us to rule out a gapless Dirac spin liquid, with a gapped $\mathbb{Z}_{2}$ liquid the most natural explanation.
@@
Anderson’s original prediction for a QSL state was the two-dimensional triangular lattice antiferromagnet. With nearest-neighbor exchange only, this system orders magnetically, but a small antiferromagnetic second-nearest-neighbor exchange $J_{2}$ theoretically stabilizes a QSL phase [4, 5, 6, 7, 8, 9, 10]. Though the existence of this phase is well-accepted theoretically (although not experimentally until now), it is not clear what kind of QSL such a state would be. Proposals include a gapless $\mathbb{U}_{1}$ Dirac QSL [6, 10, 11, 12], a valence bond crystal [13, 14], a gapped $\mathbb{Z}_{2}$ QSL [15, 4, 5, 16], or a chiral spin liquid [17]. Because numerical simulations are limited by finite size, theoretical results are ambiguous [18, 13]. The best (and perhaps only) way to resolve this question would be to find a real material which harbors the triangular lattice QSL ground state.
@@
A very promising class of materials is the Yb delafossites $A$YbSe2 where $A$ is an alkali metal [24, 25, 26, 27, 28, 29]. These form ideal triangular lattices of magnetic Yb3+, and appear to approximate the Heisenberg $J_{2}/J_{1}$ model [30], see Fig. 1. Of these, CsYbSe2 and KYbSe2 have been observed to order magnetically at zero field [28, 29]. However, following a trend in the periodic table that a smaller $A$-site element enhances $J_{2}$ and destabilizes order [30], no long-range magnetic order has been observed in NaYbSe2 [24, 27], which makes it a prime candidate for a QSL ground state. Importantly, the tunability of these compounds means that the QSL phase can be approached systematically from magnetic order (Fig. 1). This allows for greater confidence and rigor than studying a single compound in isolation would.
@@
Previous NaYbSe2 studies reported a diffuse neutron spectrum that was interpreted in terms of spinon Fermi surface excitations from a QSL [27], but because of 3% Na site disorder on those samples, it is not clear whether the magnetic order and coherent excitations were destroyed by small amounts of disorder (as in the ill-fated Yb3+ QSL candidates YbMgGaO4 [31] and Yb2Ti2O7 [32]). To further clarify NaYbSe2, we measured the inelastic neutron spectra, AC calorimetry, and AC susceptibility with high quality samples. We observe coherent excitations, lack of magnetic order, and evidence in bulk susceptibility of a 2.1 $\mu$eV gap at low temperature. This is strong evidence for a QSL ground state in NaYbSe2 and a gapped QSL on the triangular lattice.
@@
The neutron spectra at 100 mK, shown in Fig. 5, show a highly dispersive continuum of excitations with a well-defined lower bound, similar to KYbSe2 [28] (see supplemental materials for experimental details). This is qualitatively different from the spectra measured by Dai et al [27] on the 3% Yb/Na site-mixed sample which in contrast showed smeared our continua in k-space and diffuse spectra extending to low energies in many regions of reciprocal space. (Later in the text, we will explain why we believe our samples are free from mixing disorder.) Here, the only region of reciprocal space which has appreciable intensity down to low energies is $(1/3,1/3,0)$, corresponding to the 120 magnetic order seen in sister compounds KYbSe2 [28] and CsYbSe2 [29]. Down to 50 ${\rm\mu eV}$ (the limit before the incoherent scattering on the elastic line obscures the scattering energy for the incident energy of $E_{i}=1$ meV), no gap in the spectrum is resolved.
@@
For comparison we also show matrix product state (MPS) calculated spectra in 5f-i with $J_{2}/J_{1}=0.071$ (this value derived from finite field non-linear spin wave fits [30]), at varying levels of exchange anisotropy $\Delta$ (see Supplemental Materials). The boundary to the quantum spin liquid phase for the isotropic model is at $J_{2}/J_{1}=0.063$ calculated using neural quantum states (see Supplemental Materials) locating the material in the theoretically predicted QSL phase for weak anisotropies. Because of finite size lattice effects the calculated spectra are gapped, and it is difficult to make quantitative comparisons between theory and neutron experiments. Nevertheless, the calculated spectra are consistent with the observed spectra, corroborating the idea that a $J_{2}/J_{1}$ model with easy-plane anisotropy is an appropriate model for NaYbSe2.
Figure 2: Neutron scattering data on NaYbSe2. Panel a shows the triangular crystal structure and reciprocal space vectors, with $hh$ vertical. Panels b and c show neutron scattering in the $hh\ell$ scattering plane, integrated over $\ell<4.5$ reciprocal lattice units (rlu) with incident neutron energies $E_{i}=3.32$ meV and $1.55$ meV respectively. 12 K data have been subtracted as a background, see supplemental materials. Panel d shows the scattering at K$=(1/3,1/3,0)$ as a function of energy with $E_{i}=1.0$ meV and $1.55$ meV. To an energy resolution of 50 $\rm\mu$eV, the spectrum is gapless. Panel e shows the temperature-subtracted scattering at M$=(1/2,1/2,0)$ with $E_{i}=3.32$ meV, and it is unclear whether the spectrum is gapped or gapless at M. Panels f-i show MPS simulated scattering of NaYbSe2 with varying levels of anisotropy. Note the broadened signal due to finite size effects.
@@
To investigate the magnetic state to lower temperatures, we measured AC susceptibility down to 20 mK with AC and DC field applied along the $a$ and $c$ directions on NaYbSe2 (see Supplemental Materials). In this case we observe a clear magnetization plateau in the $B\parallel a$ direction at 5 T, but not for $B\parallel c$ (note these data were collected simultaneously on two separate crystals mounted on two separate susceptometers mounted on the same dilution refrigerator). This agrees with previous measurements [25], and indicates an easy-plane exchange anisotropy in NaYbSe2: in the perfectly isotropic triangular model, 1/3 magnetization plateaux appear both in-plane and out-of-plane, but the out-of-plane plateau is suppressed by planar anisotropy [34, 35, 36], although the in-plane magnetism still has a continuous rotation symmetry and similar physics is preserved.
However, the most important feature in susceptibility is the low-field drop in susceptibility at 23 mK, shown in Fig 4b and e. This drop occurs in both the $B\parallel a$ and $B\parallel c$ data. Observing such a feature at such low temperatures is prima facie evidence of high crystalline quality: any magnetic randomness or disorder in the material must involve an energy smaller than $\sim k_{B}T=2.2\>\mu$eV, or else such a feature would be suppressed. Furthermore, there is no detectable frequency dependence in either direction, shown in Fig. 4c and f, indicating that it is not a spin-freezing transition. Rather, this indicates either a magnetic ordering transition, or a gap opening in the magnetic excitations.
@@
Figure 4: Magnetic susceptibility of NaYbSe2 with $B\parallel a$ (a-c) and $B\parallel c$ (d-f). The left panels (a and d) show the field-dependence, showing a 5 T drop in susceptibility from a magnetization plateau in-plane, but no plateau out-of-plane. The center panels (b and e) show the temperature dependence at low fields, with the marked drop in susceptibility at 23 mK, normalized by the peak susceptibility value. The right panels (c and f) show the frequency dependence of the drop (where susceptibility is normalized by frequency $\nu$).
If the 23 mK susceptibility feature were a magnetic ordering transition, this would indicate that NaYbSe2 is extremely close to a QSL phase, closer than any other triangular delafossite materials [30]. The order would be extremely subtle, as the temperature $\frac{23\>{\rm mK}}{J_{1}/k_{B}}\approx 1/280$ means the amount of energy available for any type of order is very small, i.e., the order parameter saturates to a tiny value, and the system is left mostly fluctuating. However, the hypothesis of 23 mK magnetic order is unlikely, for three reasons.
First, the isotropy: the drop in susceptibility is qualitatively the same along $a$ and $c$. The only difference at zero field is a slightly higher peak temperature for $H\parallel c$ at 24 mK. Antiferromagnetic order (especially the coplanar order expected for NaYbSe2 with planar anisotropy) should produce a very different response for the field along which the spins order. Meanwhile, a gapped spectrum produces an isotropic response, consistent with what is observed here.
Second, the magnitude: the drop in susceptibility is more than 5% without background susceptibility subtracted, whereas an ordering transition at such extremely low temperatures (in comparison to a $\sim 15$ K bandwidth) would indicate very weak order with an extremely small recovered entropy, and a correspondingly weak signal in bulk properties. Indeed, in the sister compound KYbSe2, despite a clear magnetic order transition at 290 mK in heat capacity and neutron diffraction [28], the ordering feature in susceptibility is essentially invisible (see Supplemental Fig. 11). If the NaYbSe2 shows the same antiferromagnetic ordering and the ordering temperature is an order of magnitude lower in temperature than KYbSe2, we would expect a much weaker feature in susceptibility. Instead we observe a very strong feature, which is evidence for it being from a gapped spectrum. This abrupt drop is also inconsistent with spin glass behavior, where the ac susceptibility typically shows a frequency-dependent peak and symmetric decrease both above and below the peak temperature.
Third, fine-tuning: such a low transition temperature would require the system to sit exactly on the boundary between the QSL and AFM order. If we assume the system lies within the $120^{\circ}$ ordered phase, the dynamical exponent of the critical point would be $z=1$ (linearly dispersing zero modes). This implies that the effective dimension of the theory that describes the quantum phase transition is $D=d+z=3+1=4$. Since $D=4$ is the upper critical dimension (Gaussian fixed point), we expect the behavior of $T_{N}(J_{2}/J_{1})$ to be mean field like, i.e., $T_{N}\propto\sqrt{J_{2}^{c}-J_{2}}$ where $J_{2}^{c}$ is the critical value of $J_{2}$. This means that the boundary becomes vertical in $J_{2}$ vs $T$ at the lowest transition temperatures. A sharp magnetic ordering transition at 23 mK (less than 0.5% of the bandwidth) would suggest a system so finely tuned to the boundary that it is much easier to believe that the system lies within an extended gap phase.
Although the evidence points towards the susceptibility feature arising from a gap in the magnetic spectrum, an alternative explanation is nuclear-dipole ordering. We consider this unlikely because (i) only 30% of Yb nuclei are magnetic, which is below the percolation threshold (50%) for the triangular lattice, and (ii) 23 mK is quite high for nuclear dipole ordering, which is typically less than 1 mK [37]. That said, there is a noticeable nuclear Schottky anomaly in the heat capacity data in Fig. 3, which indicates some splitting the energy levels of nuclear moments. However, this does not necessarily indicate static dipolar order: the 173Yb isotope (16% natural abundance) has a nonzero electric quadrupolar moment [38] whose energy levels will be split by an ionic electric field gradient at the Yb site, producing a Schottky feature without static electronic magnetism. (Furthermore, if the ordering temperature is 23 mK, a Schottky anomaly onset at 80 mK as in Ref. [25] would be much too high.) Therefore, the most natural explanation for this feature in susceptibility is a $(2.1\pm 0.1)\rm\mu eV$ gap in the magnetic spectrum, which is estimated by fitting zero field data with $e^{-\Delta/T}$ (see Supplemental Fig. 9). This is too low to have been observed in the inelastic neutron experiment, which could not resolve features below 50 $\rm\mu eV$.
According to the generalized Lieb-Schultz-Mattis theorem, the existence of a low-energy gap in the absence of a phase transition for a translationally invariant $S=1/2$ triangular antiferromagnet implies that the ground state degeneracy must have a topological origin [39, 40]. Because these materials are known to be in close proximity to a QSL phase [30], this indicates that NaYbSe2 lies within the QSL phase. This was suggested by previous refinements of the second-nearets-neighbor exchange [30], but the observation of a spin gap is far more direct evidence.
@@
A further piece of evidence in favor of QSL physics is that the quantum critical effects seen in KYbSe2 are suppressed in NaYbSe2. More specifically, the neutron spectra in KYbSe2 show energy temperature scaling [28] due to the proximity to the quantum critical point (QCP) between the 120 and QSL phase (see Figure 1). Quantum Fisher Information is a sensitive gauge to quantum criticality [41, 42] and the elevated value of $\rm nQFI=3.4(2)$ [28] in KYbSe2 indicates the influence of the QCP. In contrast $\rm nQFI=2.3(5)$ for NaYbSe2 (see Supplementary Material) is consistent with the material being beyond the QCP (where nQFI should be a maximum) and within the QSL phase where spectral intensity is more distributed [18, 12].
The existence of a low-energy gap allows us to rule out a gapless $\mathbb{U}_{1}$ Dirac QSL [6, 10, 11], but there are at least three competing theoretical options for gapped phases on the triangular lattice: (i) a resonating valence bond (gapped $\mathbb{Z}_{2}$) liquid, (ii) a valence bond crystal (VBC) [43, 44], or (iii) a chiral QSL [45, 46]. The data we present here is insufficient to fully resolve this debate, but the strong agreement with the Schwinger boson representation of a condensed $\mathbb{Z}_{2}$ liquid in KYbSe2 [28] indicates that the gap is from $\mathbb{Z}_{2}$ topological order [15, 47]. Furthermore, the chiral QSL and valence bond crystal break discrete symmetries (time-reversal and crystalline, respectively) and hence have a finite-temperature phase transition, whereas the $\mathbb{Z}_{2}$ liquid does not. We do not observe the specific heat signature of either phase transition down to 100 mK, and no susceptibility signatures between 25 mK and 400 mK. Moreover, if the small gap were caused by a spin-Peierls instability of the U(1) Dirac spin liquid [44], the momentum distribution of the integrated intensity over $\omega$ would be expected to closely resemble that of the U(1) Dirac spin liquid state. However, dynamical variational Monte Carlo calculations for $J_{2}/J_{1}=0.07$ and $J_{2}/J_{1}=0.09$ show that the K and M points have comparable integrated spectral weights [48], which starkly contrasts with experimental observations. Thus although we cannot uniquely identify the type of QSL with the measurements described here, the simplest interpretation of our results suggests a $\mathbb{Z}_{2}$ liquid, consistent with Anderson’s original proposal [1].
In conclusion, we have used neutron spectroscopy, heat capacity, and magnetic susceptibility to investigate NaYbSe2. The susceptibility feature either indicates a phase transition or a gap. If the former, NaYbSe2 is the closest triangular delafossite material yet to a QSL; if the latter, NaYbSe2 is a gapped QSL. Details of the experiment strongly suggest a gap, which means NaYbSe2 lies within the 2D triangular lattice QSL phase with a $(2.1\pm 0.1)\>{\rm\mu eV}$ gap, making it stable against perturbations. Beyond susceptibility (i), further evidence for QSL physics are (ii) the coherent excitations observed in the neutron spectra are consistent with QSL simulated spectra, (iii) no static magnetic order is observed in specific heat down to 100 mK, and (iv) quantum entanglement witnesses indicates NaYbSe2 has less divergent intensity than KYbSe2, and is within the QSL phase. The presence of a gap allows us to rule out the gapless $\mathbb{U}_{1}$ and suggests a gapped $\mathbb{Z}_{2}$ liquid, but determining the precise nature of this ground state requires further investigation. Thus, over 50 years after Anderson’s original proposal, we finally have a clean experimental realization of a triangular lattice QSL phase.

### v2

# Spectrum and low-temperature bulk properties of triangular quantum spin liquid candidate NaYbSe2

@@ Abstract
We report neutron scattering, pressure-dependent AC calorimetry, and AC magnetic susceptibility measurements of triangular lattice NaYbSe2. We observe a continuum of scattering, which is reproduced by matrix product simulations, and no phase transition is detected in any bulk measurements. Comparison to heat capacity simulations suggest the material is within the Heisenberg spin liquid phase. AC Susceptibility shows a frequency-dependent peak at 40 mK, as has been observed in several triangular magnets.
@@
Anderson’s original prediction for a QSL state was the two-dimensional triangular lattice antiferromagnet. With nearest-neighbor exchange only, this system orders magnetically, but a small antiferromagnetic second-nearest-neighbor exchange $J_{2}$ theoretically stabilizes a QSL phase [4, 5, 6, 7, 8, 9, 10]. Though the existence of this phase is well-accepted theoretically, it is not clear what kind of QSL such a state would be. Proposals include a gapless $\mathbb{U}_{1}$ Dirac QSL [6, 10, 11, 12], a valence bond crystal [13, 14], a gapped $\mathbb{Z}_{2}$ QSL [15, 4, 5, 16], or a chiral spin liquid [17]. Because numerical simulations are limited by finite size, theoretical results are ambiguous [18, 13]. The best (and perhaps only) way to resolve this question would be to find a real material which harbors the triangular lattice QSL ground state.
@@
A very promising class of materials is the Yb delafossites AAYbSe2 where $A$ is an alkali metal [24, 25, 26, 27, 28, 29]. These form ideal triangular lattices of magnetic Yb3+, and appear to approximate the Heisenberg $J_{2}/J_{1}$ model [30], see Fig. 1. Of these, CsYbSe2 and KYbSe2 have been observed to order magnetically at zero field [28, 29]. However, following a trend in the periodic table that a smaller $A$-site element enhances $J_{2}$ and destabilizes order [30], no long-range magnetic order has been observed in NaYbSe2 [24, 27], which makes it a prime candidate for a QSL ground state. Importantly, the tunability of these compounds means that the QSL phase can be approached systematically from magnetic order (Fig. 1). This allows for greater confidence and rigor than studying a single compound in isolation would.
@@
Previous NaYbSe2 studies reported a diffuse neutron spectrum that was interpreted in terms of spinon Fermi surface excitations from a QSL [27], but because of 3% Na site disorder on those samples, it is not clear whether the magnetic order and coherent excitations were destroyed by small amounts of disorder (as in the ill-fated Yb3+ QSL candidates YbMgGaO4 [31] and Yb2Ti2O7 [32]). To further clarify NaYbSe2, we measured the inelastic neutron spectra, AC calorimetry, and AC susceptibility with high quality samples. We observe coherent excitations, lack of magnetic order, and a 40 mK frequency-dependent freezing transition. This is strong evidence for a QSL ground state in NaYbSe2 that is disrupted by dilute randomness or impurities at the lowest energies.
@@
The neutron spectra at 100 mK, shown in Fig. 2, show a highly dispersive continuum of excitations with a well-defined lower bound, similar to KYbSe2 [28] (see supplemental materials for experimental details). This is qualitatively different from the spectra measured by Dai et al [27] on the 3% Yb/Na site-mixed sample which in contrast showed smeared our continua in k-space and diffuse spectra extending to low energies in many regions of reciprocal space. (Later in the text, we will explain why we believe our samples have lower levels of mixing disorder.) Here, the only region of reciprocal space which has appreciable intensity down to low energies is $(1/3,1/3,0)$, corresponding to the 120 magnetic order seen in sister compounds KYbSe2 [28] and CsYbSe2 [29]. Down to 50 $\rm\mu eV$} (the limit before the incoherent scattering on the elastic line obscures the scattering energy for the incident energy of $E_{i}=1$ meV), no gap in the spectrum is resolved.
@@
For comparison we also show matrix product state (MPS) calculated spectra in 2f-i with $J_{2}/J_{1}=0.071$ (this value derived from finite field non-linear spin wave fits [30]), at varying levels of exchange anisotropy $\Delta$ (see Supplemental Materials). The boundary to the quantum spin liquid phase for the isotropic model is at $J_{2}/J_{1}=0.063$ calculated using neural quantum states (see Supplemental Materials) locating the material in the theoretically predicted QSL phase for weak anisotropies. Because of finite size lattice effects the calculated spectra are gapped, and it is difficult to make quantitative comparisons between theory and neutron experiments. Nevertheless, the calculated spectra are consistent with the observed spectra, corroborating the idea that a $J_{2}/J_{1}$ model with easy-plane anisotropy is an appropriate model for NaYbSe2.
[…]
@@
To investigate the magnetic state to lower temperatures, we measured AC susceptibility down to 20 mK with AC and DC field applied along the $a$ and $c$ directions on NaYbSe2, and again down to 0.7 mK in zero-field for the AC field along $a$ (see Supplemental Materials). In this case we observe a clear magnetization plateau in the $B\parallel a$ direction at 5 T, but not for $B\parallel c$ (note these data were collected simultaneously on two separate crystals mounted on two separate susceptometers mounted on the same dilution refrigerator). This agrees with previous measurements [25], and indicates an easy-plane exchange anisotropy in NaYbSe2: in the perfectly isotropic triangular model, 1/3 magnetization plateaux appear both in-plane and out-of-plane, but the out-of-plane plateau is suppressed by planar anisotropy [34, 35, 36], although the in-plane magnetism still has a continuous rotation symmetry and similar physics is preserved.
@@
Figure 4: Magnetic susceptibility of NaYbSe2 with $B\parallel a$ (a-c) and $B\parallel c$ (d-e). The left panels (a and d) show the field-dependence, showing a 5 T drop in susceptibility from a magnetization plateau in-plane, but no plateau out-of-plane. The center panels (b and e) show the temperature dependence at low fields and various frequencies, with a symmetric frequency-dependent peak at 40 mK. Panel c shows temperature dependent susceptibility measured in zero dc magnetic field down to 0.7 mK with an excitation field of 0.01 Oe, showing a broad maximum around 40 mK. Panel f shows Arrhenius analysis of the characteristic temperature $T_{f}$ defined by the peak position of the in-plane and the out-of-phase susceptibility $\chi^{\prime\prime}$. The data follow an Arrhenius behavior, $\ln(\nu)\propto 1/T_{f}$.
Susceptibility also shows a low-field peak at 40 mK as shown in Fig. 4b and e. (Originally this was misinterpreted to be a gap when measuring with a 0.9 Oe excitation field see the Supplemental Information.) The frequency-dependence in-plane and out-of-plane can be quantified by defining a characteristic temperature $T_{f}$ at the peak position of $\chi^{\prime\prime}$, which follows an activated form described by the Arrhenius relation
$\nu=\nu_{0}\exp\left(-\frac{E_{a}}{k_{B}T_{f}}\right),$ (1)
where $\nu_{0}$ is the attempt frequency and $E_{a}$ is an effective activation energy. For $B_{ac}\parallel a$, the data are well described by $\nu_{0}\approx 3\times{10}^{8}$ Hz and $E_{a}/k_{B}\approx 345$ mK, while for $B_{ac}\parallel c$ we obtain $\nu_{0}\approx 3.6\times{10}^{8}$ Hz and a slightly larger activation scale $E_{a}/k_{B}\approx 463$ mK. Moreover, fitting the data to a Vogel–Fulcher form, $\nu=\nu_{0}\exp[-E_{a}/k_{B}(T_{v}-T_{0})]$, does not yield a stable or physically meaningful Vogel–Fulcher temperature $T_{0}$, disfavoring an interpretation in terms of a cooperative glassy transition with a well-defined freezing temperature. Such slow dynamics can naturally coexist with the absence of a sharp thermodynamic anomaly and with the persistence of fast spin fluctuations, such as inelastic neutron scattering (INS), supporting a spin-liquid–proximate regime in which pronounced slowing down appears only at the longest time scales accessible to low-frequency susceptibility measurements.
This type of very low-temperature frequency dependent peak has been seen before in triangular QSL candidates [37, 38, 39, 40], but it remains an open question how to interpret. Frequency-dependent susceptibility peaks are characteristic of spin glasses [41], but the temperature $\frac{40\>{\rm mK}}{J_{1}/k_{B}}\approx 1/160$ is much lower than the energy scale of the exchange, and the degree of correlations above $T_{f}$ is very strong. Thus the ground state likely includes a significant degree of quantum entanglement, perhaps in a two-dimensional version of a random-singlet phase [40] driven by small randomness in the ground state. When the excitation field is lowered to 0.01 Oe (Fig. 4c), the peak persists, indicating it is an intrinsic zero-field feature. Below 40 mK the susceptibility continues to decrease down to 0.7 mK with no clear features.
@@
The key question about NaYbSe2 is whether it lies within the QSL phase. The data here are consistent with NaYbSe2 being within the QSL phase above $T=40$ mK, given the complete lack of magnetic order. A further piece of evidence in favor of QSL physics is that the quantum critical effects seen in KYbSe2 are suppressed in NaYbSe2. More specifically, the neutron spectra in KYbSe2 show energy temperature scaling [28] due to the proximity to the quantum critical point (QCP) between the 120 and QSL phase (see Figure 1). Quantum Fisher Information is a sensitive gauge to quantum criticality [42, 43] and the elevated value of $\rm nQFI=3.4(2)$ [28] in KYbSe2 indicates the influence of the QCP. In contrast $\rm nQFI=2.3(5)$ for NaYbSe2 (see Supplementary Material) is consistent with the material being beyond the QCP (where nQFI should be a maximum) and within the QSL phase where spectral intensity is more distributed [18, 12].
However, the presence of a low-temperature frequency-dependent peak complicates our understanding. For triangular magnets that do not magnetically order, such a feature appears to be ubiquitous (c.f. YbMgGaO4 [37], ${\mathrm{YbZn}}_{2}{\mathrm{GaO}}_{5}$ [38], and TlYbSe2 [39], and YbCuxSe2 [40]). This may indicate an inherent instability in the triangular QSL phase, as suggested by $\mathbb{U}_{1}$ Dirac simulations [44]. In this case, the weak singlet-forming distortions occur in a random pattern around dilute impurities. Then two possibilities emerge: glassy frozen regions nucleate around impurities (which are always present in real materials) or this small degree of randomness disrupts the entire phase. Resolving this question will require further experiments, but at a minimum the data here evidence a coherent QSL for $T>40$ mK—much lower temperature than the proximate QSL in KYbSe2 [28].
[…]